OpenStudy (anonymous):

On the blueprint of a house 25 millimeters represents 4 meters. The actual length of the living room is 5 meters. What is its lengthon the blueprint.

OpenStudy (lalaly):

4 meters ⇒ 25 mm 1 m ⇒ 25/4 mm 5 m ⇒ (25/4) x 5 = 125 / 4 = 31.25 mm

OpenStudy (anonymous):

Man thats rite U clearly are the best

OpenStudy (lalaly):

;)